OpenStudy (anonymous):

if f(x)=x/(x-1) and g(x)=1/(x-1) determine (f+g)(2)

OpenStudy (dumbcow):

\[f+g = \frac{x}{x-1} + \frac{1}{x-1} = \frac{x+1}{x-1}\]

jimthompson5910 (jim_thompson5910):

(f+g)(x) = f(x)+g(x) (f+g)(x) = x/(x-1) + 1/(x-1) (f+g)(2) = 2/(2-1) + 1/(2-1) (f+g)(2) = 2/1 + 1/1 (f+g)(2) = 2+1 (f+g)(2) = 3